OpenStudy (mathhelpx0):

What is the height of cylinder with a surface area of 226.08 square meters and a radius of 3 meters? (Use 3.14 for π.)

OpenStudy (anonymous):

\[SA = 2 \pi r^2 + 2 \pi r h\]

OpenStudy (anonymous):

Here as you want to find h: \[h = \frac{SA - 2 \pi r^2}{2 \pi r}\]

OpenStudy (kropot72):

Does the cylinder have closed ends or open ends?

OpenStudy (kropot72):

If the cylinder has open ends the surface area is found from: \[SA=2\times \pi \times r \times h\] \[h=\frac{SA}{2\pi r}\]